Find the equation of the line that passes the point(3,8)and has a slope of 2/3.
I need help with these. Thank you

Answer by nerdybill(6948)

Find the equation of the line that passes the point(3,8)and has a slope of 2/3.
.
Plug the given information into the "point-slope" form of:
y - y1 = m(x - x1)
y - 8 = (2/3)(x - 3)
y - 8 = (2/3)x - 2
y = (2/3)x + 6 (this is what they're looking for)
The above is in "slope-intercept" form.